What is the payment in month 204 on a 30-year loan for $925,000 at 4.85% that requires payments of $0 for the first five years and fully amortizing payments for the remaining years of the loan?
Answer #1

Payment in month 204=loan*(1+rate/12)^(12*zero payment term)*(rate/12)/(1-1/(1+rate/12)^(12*amortizing term))

=925000*(1+4.85%/12)^(12*5)*(4.85%/12)/(1-1/(1+4.85%/12)^(12*25))=6785.485734